# Toward Causal-Visual Programming: Enhancing Agentic Reasoning in Low-Code Environments*

### v1

@@ Abstract
Large language model (LLM) agents are increasingly capable of orchestrating complex tasks in low-code environments. However, these agents often exhibit hallucinations and logical inconsistencies because their inherent reasoning mechanisms rely on probabilistic associations rather than genuine causal understanding [1, 2]. This paper introduces a new programming paradigm: Causal-Visual Programming (CVP), designed to address this fundamental issue by explicitly introducing causal structures into the workflow design.
CVP allows users to define a simple “world model” for workflow modules through an intuitive low-code interface, effectively creating a Directed Acyclic Graph (DAG) that explicitly defines the causal relationships between modules [5, 6, 3]. This causal graph acts as a crucial constraint during the agent’s reasoning process, anchoring its decisions to a user-defined causal structure and significantly reducing logical errors and hallucinations by preventing reliance on spurious correlations [5, 1, 4]. To validate the effectiveness of CVP, we designed a synthetic experiment that simulates a common real-world problem: a distribution shift between the training and test environments. Our results show that a causally anchored model maintained stable accuracy in the face of this shift, whereas a purely associative baseline model that relied on probabilistic correlations experienced a significant performance drop.
@@ I-A The Rise and Challenges of LLM Agents
In recent years, intelligent agents powered by Large Language Models (LLMs) have been moving from research labs to real-world applications. These agents are particularly effective in low-code environments, where they can automate complex business processes by calling tools, accessing databases, and orchestrating UI elements [7, 9, 10]. This automation paradigm has dramatically lowered the technical barrier, enabling domain experts without programming skills to leverage the power of AI to solve problems.
@@ I-C The Solution: Causal-Visual Programming (CVP)
To address this foundational problem, this paper proposes a new programming paradigm: Causal-Visual Programming (CVP). The core idea of CVP is to make causal structure a first-class citizen in low-code environments [7, 16]. It provides an intuitive, visual interface that allows the user—a domain expert—to explicitly define the causal relationships among the operational modules in a workflow [1, 8]. This user-defined causal graph serves as a lightweight “world model”, a minimal set of assumptions that encodes human understanding of how the task environment works [3].
@@ II-B Large Language Models and Causal Reasoning
  • LLM’s Intrinsic Limitations: LLMs are trained on massive text corpora to predict the next token. This makes them skilled at capturing statistical patterns but unable to build a mechanistic causal model of how things work [1, yang2025causality, 6]. This leads to poor performance on complex tasks that require robust causal reasoning, such as economic analysis or healthcare [1, 12]. An important issue is that LLMs are highly sensitive to how causal graphs are encoded, and even powerful models like GPT-4 and Gemini-1.5 exhibit significant biases when processing them [17, 4].
@@ II-C Low-Code and Visual Programming
Low-code and visual programming platforms simplify complex software development with graphical interfaces, making them key tools for building LLM agent workflows [7].
  • Low-Code and LLM Agents: Existing low-code/no-code tools help users design and build complex task flows through a visual interface [7, 22]. Users can drag and drop modules and connect them to define the execution order, which provides more controllable and stable responses from LLM agents. These tools abstract complex prompt engineering into visual workflows, greatly enhancing the human-LLM interaction experience [7, 16].
@@ IV-A Experimental Setup
The experimental setup uses two prediction functions:
@@ V-A Mitigating Hallucinations and Improving Explainability
The CVP framework’s value extends to fundamentally addressing hallucinations and explainability. By providing a user-defined causal graph, CVP reduces information uncertainty, compelling the agent to follow a defined causal path rather than generating fabricated information [6]. This also makes the agent’s decisions transparent and traceable, simplifying the debugging process [3].
@@ VI Conclusion
In this paper, we proposed a new programming paradigm, Causal-Visual Programming (CVP), to address the issue of hallucinations and logical errors in LLM agents caused by their lack of causal understanding. CVP utilizes a human-defined causal graph as a “world model” and reasoning constraint, enabling agents to move beyond reliance on spurious correlations. A synthetic experiment quantitatively demonstrated CVP’s effectiveness, showing that a causally anchored model achieved significantly greater robustness to distribution shifts compared to an associative baseline. This work provides a novel, human-centric solution for building more trustworthy, interpretable, and generalizable AI agents.
Acknowledgment
The authors would like to thank the University of California, Irvine for providing the computational resources and research environment that made this work possible. We also acknowledge the valuable contributions and collaborative efforts of all co-authors from their respective institutions.

### v2

@@ Abstract
Large language model (LLM) agents are increasingly capable of orchestrating complex tasks in low-code environments. However, these agents often exhibit hallucinations and logical inconsistencies because their inherent reasoning mechanisms rely on probabilistic associations rather than genuine causal understanding [1, 2, 3]. This paper introduces a new programming paradigm: Causal-Visual Programming (CVP), designed to address this fundamental issue by explicitly introducing causal structures into the workflow design.
CVP allows users to define a simple world model for workflow modules through an intuitive low-code interface, effectively creating a Directed Acyclic Graph (DAG) that explicitly defines the causal relationships between modules [6, 7, 4]. This causal graph acts as a crucial constraint during the agent’s reasoning process, anchoring its decisions to a user-defined causal structure and significantly reducing logical errors and hallucinations by preventing reliance on spurious correlations [6, 2, 5]. To validate the effectiveness of CVP, we designed a synthetic experiment that simulates a common real-world problem: a distribution shift between the training and test environments. Our results show that a causally anchored model maintained stable accuracy in the face of this shift, whereas a purely associative baseline model that relied on probabilistic correlations experienced a significant performance drop.
@@ I-A The Rise and Challenges of LLM Agents
In recent years, intelligent agents powered by Large Language Models (LLMs) have been moving from research labs to real-world applications. These agents are particularly effective in low-code environments, where they can automate complex business processes by calling tools, accessing databases, and orchestrating UI elements [12, 8, 10, 11]. This automation paradigm has dramatically lowered the technical barrier, enabling domain experts without programming skills to leverage the power of AI to solve problems.
@@ I-C The Solution: Causal-Visual Programming (CVP)
To address this foundational problem, this paper proposes a new programming paradigm: Causal-Visual Programming (CVP). The core idea of CVP is to make causal structure a first-class citizen in low-code environments [16, 8, 26]. It provides an intuitive, visual interface that allows the user—a domain expert—to explicitly define the causal relationships among the operational modules in a workflow [17, 2, 9]. This user-defined causal graph serves as a lightweight “world model”, a minimal set of assumptions that encodes human understanding of how the task environment works [4].
@@ II-B Large Language Models and Causal Reasoning
  • LLM’s Intrinsic Limitations: LLMs are trained on massive text corpora to predict the next token. This makes them skilled at capturing statistical patterns but unable to build a mechanistic causal model of how things work [2, yang2025causality, 7]. This leads to poor performance on complex tasks that require robust causal reasoning, such as economic analysis or healthcare [2, 14]. An important issue is that LLMs are highly sensitive to how causal graphs are encoded, and even powerful models like GPT-4 and Gemini-1.5 exhibit significant biases when processing them [27, 5].
@@ II-C Low-Code and Visual Programming
Low-code and visual programming platforms simplify complex software development with graphical interfaces, making them key tools for building LLM agent workflows [19, 8].
  • Low-Code and LLM Agents: Existing low-code/no-code tools help users design and build complex task flows through a visual interface [8, 32]. Users can drag and drop modules and connect them to define the execution order, which provides more controllable and stable responses from LLM agents. These tools abstract complex prompt engineering into visual workflows, greatly enhancing the human-LLM interaction experience [20, 8, 26].
@@ IV-A Experimental Setup
The experimental setup uses two prediction functions to evaluate robustness under distribution shift [18]:
@@ V-A Mitigating Hallucinations and Improving Explainability
The CVP framework’s value extends to fundamentally addressing hallucinations and explainability. By providing a user-defined causal graph, CVP reduces information uncertainty, compelling the agent to follow a defined causal path rather than generating fabricated information [7]. This also makes the agent’s decisions transparent and traceable, simplifying the debugging process [21, 22, 4].
@@ VI Conclusion
In this paper, we proposed a new programming paradigm, Causal-Visual Programming (CVP), to address the issue of hallucinations and logical errors in LLM agents caused by their lack of causal understanding. CVP utilizes a human-defined causal graph as a “world model” and reasoning constraint, enabling agents to move beyond reliance on spurious correlations. A synthetic experiment quantitatively demonstrated CVP’s effectiveness, showing that a causally anchored model achieved significantly greater robustness to distribution shifts compared to an associative baseline. This work provides a novel, human-centric solution for building more trustworthy, interpretable, and generalizable AI agents. Future advances in context utilization techniques [23] will further enhance the effectiveness of CVP by enabling agents to better leverage causal graph information and maintain consistent reasoning across extended interactions.